# Geometric spectral theory of quantum graphs

### v2

@@
Geometric spectral theory
of quantum graphs
James Kennedy
Heat kernels and spectral geometry:
from manifolds to graphs
Preface
@@
These notes centre around how the spectrum of a quantum graph (that is, a Schrödinger-type differential operator defined on a metric graph) depends on the geometric and topological properties of the graph. However, the notes are aimed at non-specialists with no prior exposure to quantum graphs; in principle no more than a solid grounding in analysis and functional analysis is needed, although some exposure to the modern theory of (linear) partial differential equations is highly advantageous.
As such, the first two chapters are introductory in nature, covering the basics of weak solution theory and the spectral theorem for the Laplacian (as one might see in an advanced master’s course) and a construction of metric graphs and differential operators on them, respectively. As such, few to no citations have been included on the understanding that this material is standard/well known/folklore, although no claim is made to originality or authorship of any part (the particular organisation and presentation of the topics is my own, but unlikely original). Much of Chapter Geometric spectral theory of quantum graphs is likely similar in content to parts of [15], which was suggested background reading for the participants, although [5] is probably closer to my own way of thinking. Likewise, Chapter 1 was (with a few, marked, exceptions) not written following any particular source, but was likely influenced by a number of works (including ones partly due to me). Possible influences include, but are by no means limited to, [6, 10, 11, 12, 14, 29, 40].
The core of the matter is Chapter 2, which covers the evolution of eigenvalue estimates for quantum graph Laplacians over the last 35 years (including a brief apology for the study of geometric spectral theory in Section 1, as well as references to related topics such as Weyl asymptotics or the study of nodal domains and nodal counts). This draws far more heavily on recent literature and contains no new research, but I believe the synthesis as such is original. For the current state of the art, see, for example, [12, 13].
@@
There are a few deliberate changes from the original version to improve readability and usefulness as a standalone reference, roughly in keeping with the changes introduced for the agregação version; in particular, a few previously “staccato” passages of text and motivations have been filled out, and a bit more historical context is given in several places. More significantly, I have also yielded to the temptation to introduce a few more recent and advanced results in whose creation I was involved, which will inevitably be in more of a “survery article” style than a “lecture notes” style. To quarantine any damage to the balance and flow of the lecture notes these are collected in a new final Section 4 not present in the original version.
@@
James Kennedy
Lisbon, December 2023
Chapter 0 Laplacians on domains
1 From forms to operators
@@
1.4 Theorem (Spectral theorem).
@@
1.5 Theorem (Courant–Fischer/min-max and max-min characterisation of the eigenvalues).
@@
$\lambda_{k}$ $=\min_{u\perp u_{1},\ldots,u_{k-1}}\frac{a(u,u)}{\|u\|_{H}^{2}}=\min_{\begin{subarray}{c}U\subset V\operatorname{subspace}\\ \dim U=k\end{subarray}}\,\,\max_{0\neq u\in U}\frac{a(u,u)}{\|u\|_{H}^{2}}$ (1d)
$=\max_{\begin{subarray}{c}W\subset V\operatorname{subspace}\\ \operatorname{codim}W=k-1\end{subarray}}\,\,\max_{0\neq u\in W^{\perp}}\frac{a(u,u)}{\|u\|_{H}^{2}}.$ (1g)
Equality is attained in (1d) if $U=\operatorname{span}\{u_{1},\ldots,u_{k}\}$ and in (1g) if $W=(\operatorname{span}\{u_{1},\ldots,u_{k-1}\})^{\perp}$.
@@
2 The spectral theorem for the Laplacian
@@
which will play the role of the space $H$ from Section 1.
To define our $V$, we need to introduce Sobolev spaces, i.e. spaces of weakly differentiable functions. (At an intuitive level it is often sufficient to imagine these functions as being piecewise-smooth functions (which may have “kinks”), and to know that all basic theorems from vector calculus continue to hold for them, and that they form a complete space with respect to the natural norm of the space. In particular, we can apply our abstract theorems to them.)
2.1 Definition.
@@
Higher order weak partial derivatives are defined accordingly: for any multiindex $\alpha=(\alpha_{1},\ldots,\alpha_{d})\in\mathbb{N}_{0}^{d}$, of order $|\alpha|=\alpha_{1}+\ldots+\alpha_{d}$, $\frac{\partial^{\alpha}u}{\partial x^{\alpha}}=\frac{\partial^{|\alpha|}u}{\partial x_{1}^{\alpha_{1}}\ldots\partial x_{d}^{\alpha_{d}}}$ is defined to be the locally integrable function $g_{\alpha}$, if one exists, such that
@@
2.2 Definition (Sobolev spaces).
@@
$W^{k,p}(\Omega)=\{u\in L^{p}(\Omega):\text{all weak derivatives of $u$ up to order $k$ exist and belong to $L^{p}(\Omega)$}\}$
@@
to be the subspace of $W^{k,p}(\Omega)$ consisting of all functions which can be approximated in the $W^{k,p}$-norm by smooth functions of compact support. When $p=2$, write $H^{k}(\Omega)$ in place of $W^{k,2}(\Omega)$, and $H^{k}_{0}(\Omega)$ in place of $W^{k,2}_{0}(\Omega)$.
@@
for all $u,v\in H^{1}(\Omega)$ (or $H^{1}_{0}(\Omega)$). In the case of $H^{1}_{0}(\Omega)$, the alternative expression $\int_{\Omega}\nabla u\cdot\overline{\nabla v}\,\textrm{d}x$ already defines an inner product, as follows from Friedrichs’ inequality. We will not need this.
@@
2.6 Theorem (Embedding theorems, small selection).
@@
To define Laplace-type operators on $\Omega$ (plus suitable boundary conditions on $\partial\Omega$), we use the form approach of Section 1: we define a form
@@
2.9 Theorem (Spectral theorem for the Dirichlet and Neumann Laplacians).
@@
$-\Delta^{D}:\qquad$ $\lambda_{1}\leq\lambda_{2}\leq\lambda_{3}\leq\ldots\to\infty,$
$-\Delta^{N}:\qquad$ $\mu_{1}\leq\mu_{2}\leq\mu_{3}\leq\ldots\to\infty.$
@@
[…]
@@
3 Application: the heat equation
@@
Loosely, $\lambda_{1}$ and $\mu_{2}$ determine the rate of convergence/heat loss from $\Omega$. In the Dirichlet case, where there is “perfect cooling” at the boundary since $\lambda_{1}>0$,
@@
exponentially as $t\to\infty$, with exponent (at most) $\lambda_{1}$. In the Neumann case, which represents perfect insulation, i.e. no heat loss through the boundary, since $\mu_{1}=0$ and $\mu_{2}>0$,
@@
Chapter 1 Quantum graphs
1 Metric graphs
@@
This distinction is more directly visible in the world of discrete graphs; for example, the adjacency matrix of the graph is symmetric if the graph is undirected and not generally symmetric if it is directed. (Another example: Twitter, with its network of followers, defines a directed discrete graph whose vertices are the users; Facebook, with its network of “friends”, defines an undirected discrete graph.) Our metric graphs will be undirected; in other words, the choice of the orientation of the intervals will not matter.
@@
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}0_{1}}$${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}1_{1}}$${\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}0_{2}}$${\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}1_{2}}$${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}0_{3}}$${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}1_{3}}$$e_{1}$$e_{2}$$e_{3}$${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}0_{1}}$${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}1_{1}}$${\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}0_{2}}$${\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}1_{2}}$${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}1_{3}}$${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}0_{3}}$$e_{1}$$e_{2}$$e_{3}$
@@
Proof.
Tedious exercise. ∎
A metric graph $\Gamma$ consisting of finitely many edges of finite length is thus generally called a compact metric graph in the literature, cf. [14, Definition 1.3.4]. Finally, we emphasise that we will generally treat $\Gamma$ as a metric space of points $x\in\Gamma$ rather than the pair $(V,E)$ of a vertex set and an edge set (although whenever convenient we will identify the two).
2 Function spaces
@@
[…]
@@
[…]
@@
2.4 Lemma.
$f\in C(\Gamma)$ iff $f$ is continuous with respect to the metric/topology of Section 1; in particular, if $v\sim e$ and $x_{n}\in e$ with $x_{n}\to v$ in $\Gamma$, then $f(x_{n})\to f(v)$.
2.5 Theorem.
Under our assumptions on $\Gamma$ (i.e., for $\Gamma$ a compact metric graph), the quantity
@@
2.6 Lemma (Embedding theorem).
@@
We may thus characterise $W^{1,p}$-functions as continuous functions (up to the choice of the correct representative) which have a weak derivative in $L^{p}$; in particular, $W^{1,p}$-functions have well-defined values at all points. This is a one-dimensional equivalent of the trace operator introduced in Section 2.
2.7 Definition.
@@
[…]
2.9 Remark.
@@
2.10 Remark (Dummy vertices).
@@
2.12 Remark.
It is easy to define higher order derivatives edgewise: for example, if $f\in W^{1,p}(\Gamma)$, and $e\in E$, we may write $f|_{e}^{\prime\prime}\in L^{p}(e)$, and so $f|_{e}\in W^{2,p}(e)$, if there is a function $g\in L^{p}(e)$ (which we will call $f|_{e}^{\prime\prime}$) such that
@@
for all $\phi\in C_{c}^{\infty}(e)$, i.e., for all functions $\phi$ which are supported in the interior of the edge $e$. (Compare with Definition 2.4.)
@@
3 Laplacian and Schrödinger operators on metric graphs

### v3

@@
Abstract
These are lecture notes from a course given at the summer school “Heat kernels and spectral geometry: from manifolds to graphs” in Bregenz, Austria, 2022. They are designed to be accessible to doctoral level students, and include background chapters on Laplacians on domains and quantum graphs before moving on to specialised topics involving the dependence and optimisation of operator eigenvalues on a metric graph in function of the graph geometry, drawn in part from the recent literature.
keywords:
Laplacian, metric graph, spectrum, eigenvalue estimates.
@@ Abstract
[…]
1 Introduction
@@
These notes centre around how the spectrum of a quantum graph (i.e. a Schrödinger-type differential operator defined on a metric graph) depends on the geometric and topological properties of the graph. However, the notes are aimed at non-specialists with no prior exposure to quantum graphs; in principle no more than a solid grounding in analysis and functional analysis is needed, although some exposure to the modern theory of (linear) partial differential equations is highly advantageous.
As such, the first two chapters are introductory in nature, covering the basics of weak solution theory and the spectral theorem for the Laplacian (as one might see in an advanced master’s course) and a construction of metric graphs and differential operators on them, respectively. As such, few to no citations have been included on the understanding that this material is standard/well known/folklore, although no claim is made to originality or authorship of any part (the particular organisation and presentation of the topics is my own, but unlikely original). Much of Chapter 2 is likely similar in content to parts of [16], which was suggested background reading for the participants, although [5] is probably closer to my own way of thinking. Likewise, Chapter 3 was (with a few, marked, exceptions) not written following any particular source, but was likely influenced by a number of works (including ones partly due to me). Possible influences include, but are by no means limited to, [6, 11, 12, 13, 15, 30, 41].
The core of the matter is Chapter 4, which covers the evolution of eigenvalue estimates for quantum graph Laplacians over the last 35 years (including a brief apology for the study of geometric spectral theory in Section 4.1, as well as references to related topics such as Weyl asymptotics or the study of nodal domains and nodal counts). This draws far more heavily on recent literature and contains no new research, but I believe the synthesis as such is original. For the current state of the art, see, for example, [13, 14].
@@
There are a few deliberate changes from the original version to improve readability and usefulness as a standalone reference, roughly in keeping with the changes introduced for the agregação version; in particular, a few previously “staccato” passages of text and motivations have been filled out, and a bit more historical context is given in several places. More significantly, I have also introduced a few more recent and advanced results in whose creation I was involved, which will inevitably be in more of a “survey article” style than a “lecture notes” style. To quarantine any effect on the balance and flow of the lecture notes these are collected in a new final Section 4.4 not present in the original version. On the other hand, the exercises originally included for the benefit of the participants have been kept, and are in an appendix to this version.
@@
2 Laplacians on domains
2.1 From forms to operators
@@
Theorem 2.3 (Spectral theorem).
@@
Theorem 2.4 (Courant–Fischer/min-max and max-min characterisation of the eigenvalues).
@@
$\lambda_{k}$ $=\min_{u\perp u_{1},\ldots,u_{k-1}}\frac{a(u,u)}{\|u\|_{H}^{2}}=\min_{\begin{subarray}{c}U\subset V\operatorname{subspace}\\ \dim U=k\end{subarray}}\,\,\max_{0\neq u\in U}\frac{a(u,u)}{\|u\|_{H}^{2}}$ (1b)
$=\max_{\begin{subarray}{c}W\subset V\operatorname{subspace}\\ \operatorname{codim}W=k-1\end{subarray}}\,\,\max_{0\neq u\in W^{\perp}}\frac{a(u,u)}{\|u\|_{H}^{2}}.$ (1c)
Equality is attained in equation (1b) if $U=\operatorname{span}\{u_{1},\ldots,u_{k}\}$ and in equation (1c) if $W=(\operatorname{span}\{u_{1},\ldots,u_{k-1}\})^{\perp}$.
@@
2.2 The spectral theorem for the Laplacian
@@
which will play the role of the space $H$ from Section 2.1.
To define our $V$, we need to introduce Sobolev spaces, i.e. spaces of weakly differentiable functions. (At an intuitive level it is often sufficient to imagine these functions as being piecewise-smooth functions, which may have “kinks”, and to know that (1) all basic theorems from vector calculus continue to hold for them, and (2) that they form a complete space with respect to the natural norm of the space. In particular, we can apply the spectral theorem to suitable operators defined on them.)
Definition 2.5.
@@
Higher order weak partial derivatives are defined accordingly: for any multi-index $\alpha=~{}(\alpha_{1},\ldots,\alpha_{d})\in\mathbb{N}_{0}^{d}$, of order $|\alpha|=\alpha_{1}+\ldots+\alpha_{d}$, $\frac{\partial^{\alpha}u}{\partial x^{\alpha}}=\frac{\partial^{|\alpha|}u}{\partial x_{1}^{\alpha_{1}}\ldots\partial x_{d}^{\alpha_{d}}}$ is defined to be the locally integrable function $g_{\alpha}$, if one exists, such that
@@
Definition 2.6 (Sobolev spaces).
@@
$W^{k,p}(\Omega)=\{u\in L^{p}(\Omega):\text{all weak derivatives of $u$ up to order $k$ exist and are in $L^{p}(\Omega)$}\}$
@@
to be the subspace of $W^{k,p}(\Omega)$ consisting of all functions which can be approximated in the $W^{k,p}$-norm by smooth functions of compact support. When $p=2$, we write $H^{k}(\Omega)$ in place of $W^{k,2}(\Omega)$, and $H^{k}_{0}(\Omega)$ in place of $W^{k,2}_{0}(\Omega)$.
@@
for all $u,v\in H^{1}(\Omega)$ (or $H^{1}_{0}(\Omega)$). Note that in the case of $H^{1}_{0}(\Omega)$, the alternative expression $\int_{\Omega}\nabla u\cdot\overline{\nabla v}\,\textrm{d}x$ already defines an inner product, as follows from Friedrichs’ inequality. We will not need this.
@@
Theorem 2.10 (Embedding theorems, small selection).
@@
To define Laplace-type operators on $\Omega$ (plus suitable boundary conditions on $\partial\Omega$), we use the form approach of Section 2.1: we define a form
@@
Theorem 2.13 (Spectral theorem for the Dirichlet and Neumann Laplacians).
@@
$-\Delta^{D}:$ $\lambda_{1}\leq\lambda_{2}\leq\lambda_{3}\leq\ldots\to\infty,$
$-\Delta^{N}:$ $\mu_{1}\leq\mu_{2}\leq\mu_{3}\leq\ldots\to\infty.$
@@
$\frac{a(u,u)}{\|u\|_{2}^{2}}=\frac{\int_{\Omega}|\nabla u|^{2}\,\textrm{d}x}{\int_{\Omega}|u|^{2}\,\textrm{d}x}.$
@@
2.3 Application: the heat equation
@@
Loosely, $\lambda_{1}$ and $\mu_{2}$ determine the rate of convergence/heat loss from $\Omega$. In the Dirichlet case, where there is “perfect cooling” at the boundary, since $\lambda_{1}>0$,
@@
exponentially as $t\to\infty$, with exponent (at least) $\lambda_{1}$. In the Neumann case, which represents perfect insulation, i.e. no heat loss through the boundary, since $\mu_{1}=0$ and $\mu_{2}>0$,
@@
3 Quantum graphs
3.1 Metric graphs
@@
This distinction is more directly visible in the world of discrete graphs; for example, the adjacency matrix of the graph is symmetric if the graph is undirected and not generally symmetric if it is directed. (Another example: X, formerly Twitter, with its network of followers, defines a directed discrete graph whose vertices are the users; Facebook, with its network of “friends”, defines an undirected discrete graph.) Our metric graphs will be undirected; in other words, the choice of the orientation of the intervals will not matter.
@@
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0_{1}}$${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1_{1}}$${\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}0_{2}}$${\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}1_{2}}$${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0_{3}}$${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1_{3}}$$e_{1}$$e_{2}$$e_{3}$${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0_{1}}$${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1_{1}}$${\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}0_{2}}$${\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}1_{2}}$${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1_{3}}$${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0_{3}}$$e_{1}$$e_{2}$$e_{3}$
@@
Proof 3.7.
Tedious exercise.
A metric graph $\Gamma$ consisting of finitely many edges of finite length is thus generally called a compact metric graph in the literature, cf. [15, Definition 1.3.4]. Finally, we emphasise that we will generally treat $\Gamma$ as a metric space of points $x\in\Gamma$ rather than the pair $(V,E)$ of a vertex set and an edge set (although whenever convenient we will identify the two).
3.2 Function spaces
@@
$\|f\|_{p}=\begin{cases}\left(\sum_{k=1}^{m}\int_{e_{k}}|f(x)|^{p}\,\textrm{d}x\right)^{1/p},\qquad&1\leq p<\infty,\\ \max_{k=1,\ldots,m}\|f|_{e_{k}}\|_{L^{\infty}(e_{k})}=\operatorname*{ess\,sup}_{x\in\Gamma}|f(x)|,\qquad&p=\infty,\end{cases}$ (10)
@@
$\langle f,g\rangle=\int_{\Gamma}f\overline{g}\,\textrm{d}x\equiv\sum_{k=1}^{m}\int_{e_{k}}f(x)\overline{g(x)}\,\textrm{d}x.$
@@
Lemma 3.11.
$f\in C(\Gamma)$ iff $f$ is continuous with respect to the metric/topology of Section 3.1; in particular, if $v\sim e$ and $x_{n}\in e$ with $x_{n}\to v$ in $\Gamma$, then $f(x_{n})\to f(v)$.
Theorem 3.12.
Under our assumptions on $\Gamma$ (i.e. for $\Gamma$ a compact metric graph), the quantity
@@
Lemma 3.13 (Embedding theorem).
@@
We may thus characterise $W^{1,p}$-functions as continuous functions (up to the choice of the correct representative) which have a weak derivative in $L^{p}$; in particular, $W^{1,p}$-functions have well-defined values at all points. This is a one-dimensional equivalent of the trace operator introduced in Section 2.2.
Definition 3.14.
@@
$\langle f,g\rangle=\int_{\Gamma}f^{\prime}\overline{g}^{\prime}+f\overline{g}\,\textrm{d}x\equiv\sum_{k=1}^{m}\int_{e_{k}}f^{\prime}(x)\overline{g^{\prime}(x)}+f(x)\overline{g(x)}\,\textrm{d}x.$
Remark 3.16.
@@
Remark 3.17 (Dummy vertices).
@@
Remark 3.19.
It is easy to define higher order derivatives edgewise: for example, if $f\in\nolinebreak W^{1,p}(\Gamma)$, and $e\in E$, we may write $f|_{e}^{\prime\prime}\in L^{p}(e)$, and so $f|_{e}\in W^{2,p}(e)$, if there is a function $g\in L^{p}(e)$ (which we will call $f|_{e}^{\prime\prime}$) such that
@@
for all $\phi\in C_{c}^{\infty}(e)$, i.e. for all functions $\phi$ which are supported in the interior of the edge $e$. (Compare with Definition 2.8.)
@@
3.3 Laplacian and Schrödinger operators on metric graphs